Three heating elements are connected in parallel. What is the quickest action you can take to measure the resistance of just one element?

a. measure the resistance across all three heaters and multiply by three
b. measure the resistance of all of the heaters and divide by three
c. disconnect one side of the heater from the circuit
d. remove both connections to the heater and measure in parallel
Answer #1

B. Measure all resistances of heaters and decided by 3

3 Resistances in parallel then equilent is

1/Req= 1/R+1/R+1/R

1/Req=3/R

Req=R/3